# ComFace: Facial Representation Learning with Synthetic Data for Comparing Faces

### v1

@@ 4.7 Visualization
We calculate saliency maps to provide the interpretability of models. The saliency maps are obtained from the final convolutional layer of the pre-trained backbones via Eigen-CAM [43], which visualizes the principle components of learned representations without relying on class relevance scores. Figure 4 illustrates the saliency maps of four models for a face image from DISFA [42, 41]. Models using ImageNet and SimCLR focus on regions other than the face (e.g., clothing and hair), which is expected due to the fact that the models are trained on general images (i.e., ImageNet). The model using VGGFace2 focuses on the entire face since it is trained for face recognition. Our ComFace focuses on the center of the face, which includes AU6, AU12, and eyelid and nose shapes that are affected by edema [2] (see the original image in Fig. 4). We believe that ComFace captures the center of the face, where intra-personal facial changes often appear, resulting in high transfer performance of downstream tasks for comparing faces.
5 Conclusion
This paper introduces facial representation learning (FRL) using synthetic images for comparing faces. We obtain the following answers from the experimental results: (i) ComFace successfully acquires representations that capture intra-personal facial change. (ii) Our transfer performance in representation learning using synthetic images is comparable to or better than SoTA representation learning methods using real images. (iii) Our approach of comparing two face images generalizes well to new patients and environmental conditions.
Limitations: Our FRL with synthetic images is adapted only to downstream tasks for comparing faces. Future work should adapt FRL using synthetic images to a variety of face analysis tasks. We should also validate generative models for face synthesis other than StyleGANs (e.g., diffusion models).
While we acknowledge the above limitations, we hope this work opens new possibilities for representation learning with synthetic data in the face domain.

### v3

@@ 4.6 Visualization
We calculate saliency maps to provide the interpretability of models. The saliency maps are obtained from the final block of the pre-trained backbones via Eigen-CAM [44], which visualizes the principle components of learned representations without relying on class relevance scores. Figure 4 illustrates the saliency maps of four models for a face image from DISFA [43, 42]. Models using ImageNet and SimCLR focus on regions other than the face (e.g., clothing and hair), which is expected due to the fact that the models are trained on general images (i.e., ImageNet). The model using VGGFace2 focuses on the entire face since it is trained for face recognition. Our ComFace focuses on the center of the face, which includes AU6, AU12, and eyelid and nose shapes that are affected by edema [2] (see the original image in Fig. 4). We believe that ComFace captures the center of the face, where intra-personal facial changes often appear, resulting in high transfer performance of downstream tasks for comparing faces.
5 Conclusion and Societal Impact
This paper introduces facial representation learning (FRL) using synthetic images for comparing faces. We obtain the following answers from the experimental results: (i) ComFace successfully acquires representations that capture intra-personal facial change. (ii) Our transfer performance in representation learning using synthetic images is comparable to or better than SoTA representation learning methods using real images. (iii) Our approach of comparing two face images generalizes well to new patients and environmental conditions. In future work, we plan to explore more downstream tasks related to facial changes other than facial expression, weight, and age.
Potential Negative Societal Impact: Machine learning models using face images may include biases such as ethnicity, age, and gender. Leveraging synthetic images has the potential to reduce those biases by manipulating the generative model. However, it is important to acknowledge that our method relies on generative models trained on images crawled from a large-scale website (Flicker). The generative model inherits the biases of Flicker, and the synthetic images may contain social biases and errors.